OpenStudy (anonymous):

2t + 8 > or = to -4(t +1) t > or = to ?

OpenStudy (anonymous):

a. -7/6 b. -2 c. -6 d. 2

OpenStudy (texaschic101):

2t + 8 >= -4(t + 1) -- distribute through the parenthesis 2t + 8 >= -4t - 4 -- add 4t to both sides 2t + 4t + 8 >= -4 -- subtract 8 from both sides 6t >= -4 - 8 6t >= - 12 t >= -2